# Spineless Traversal for Layout Invalidation

### v4

@@ 3.1. Dirty bits
Incremental layout is conceptually straightforward. Each layout field has a corresponding dirty bit, which defines whether that layout field needs to be recomputed. [^5] Several layout fields can all have the same corresponding dirty bit, which is set if any of those fields need to be recomputed. When the application code writes to an attribute or property, or adds or removes layout nodes, it sets the dirty bits for all layout fields that read from those fields or node pointers. For example, in the paragraph layout of Figure 5, when a node’s height attribute is changed, its $H$ field must be marked dirty. When a subtree is deleted, its next sibling’s $HA$ field must be marked dirty. If the subtree was the last child of its parent, the parent’s $H$ field must also be marked dirty. This set of fields can be statically determined in our DSL and in fact our Megatron compiler can synthesize this marking code.
Incremental layout then has the task of clearing the dirty bits by recomputing dirty fields. Conceptually this is simple: find all dirty bits on all nodes and recompute their fields. However, when a layout field is recomputed, its value might change, and any layout fields that read the old value are now out of date. Thus, when incremental layout modifies a layout field, it must dirty all layout fields that depend on it, “propagating” dirty bits through the document. For example, in the paragraph layout of Figure 5, when a node’s $H$ field is recomputed, its $HA$ field must be marked dirty. When that $HA$ field is recomputed, its next sibling’s $HA$ field is then marked dirty in turn. Because of dirty bit propagation, the set of layout fields currently marked dirty is distinct from the set of all layout fields that incremental layout must ultimately recompute.
This second set is discovered in the process of performing incremental layout, since we only know if a layout field’s value changes when we recompute it and can compare the old and new value. For example, suppose the user is typing out a paragraph of text, laid out using Figure 5’s layout algorithm. Most edits just add text to a single line, not affecting its height or available width; these only end up dirtying a single field on a single node. This is why incremental layout works: many interactions—especially latency-critical ones like hovers, drags, animations, and text editing—do not change fields on containing elements, and end up affecting only a small portion of the page.
@@ 3.2. Recomputation Order
Incremental layout strives to minimize the number of layout fields recomputed. As one part of this, it must ensure that, once recomputed, a layout field is not marked dirty again during that layout. Luckily, there is a simple way to guarantee this: a correct from-scratch layout always processes dependencies before the field itself, so recomputing layout fields in the same relative order ensures that each layout field is marked dirty at most once. A naive incremental layout algorithm might thus execute the same layout schedule as a from-scratch layout but skip re-computing any fields that aren’t dirty. This algorithm recomputes the minimal number of layout fields, clears all the dirty bits, and avoids ever recomputing a layout field more than once.
However, this naive algorithm isn’t particularly fast: it must access every layout node to check its dirty bits, which incurs a lot of cache misses. In the common case, most dirty bits aren’t set; these “auxiliary” accesses cause needless cache misses and end up wasting a significant fraction of run time. For incremental layout to be fast, auxiliary accesses—accesses that don’t result in field recomputations—must be minimized. We note before continuing that, while this is a quite specific and specialized problem, it is a major source of latency in existing web browsers, which in turn are both critical application platforms and also already highly optimized, meaning remaining sources of latency are particularly challenging.
@@ 3.3. The Double Dirty Bit Algorithm
The state-of-the-art Double Dirty Bit algorithm, uses summary bits to reduce auxiliary accesses; it is used, with variations, in all major rendering engines. For each dirty bit, we add a second “summary bit”, which is set for a node when that dirty bit is set anywhere in the subtree rooted at that node. Every time a dirty bit is set, the associated summary bit is set on that node and all its ancestors. When performing incremental layout, subtrees whose summary bits are clear can be skipped. Figure 7 contains pseudo-code for Double Dirty Bit as an iterator over dirty nodes.
@@
    self.dirty = True
@@
        self.dirty = False
@@
Skipping subtrees with no dirty bits reduces the number of auxiliary accesses and greatly improves performance (Garseil, [n. d.]; Panchekha and Harrelson, 2024). However, it still has some auxiliary accesses. Specifically, when a dirty bit is set on some node, summary bits will be set for all ancestors of that node, which Double Dirty Bit will have to check. Moreover, for each ancestor, Double Dirty Bit will recurse into its children, adding even more auxiliary accesses. Since layout trees are poorly balanced and often very wide, the number of auxiliary accesses can be large. For example, a single dirtied node in Figure 1 has 66 auxiliary nodes, and on larger pages the dirty-to-auxiliary ratio can be much worse.
Developers can sometimes reduce the number of auxiliary accesses by reorganizing large, complex pages, and performance monitoring tool like Google’s Lighthous (Team, [n. d.]) suggest doing so. But this can require global changes to the shape of the layout tree, which (in modern frameworks like React) requires refactoring the application as a whole. Naturally, an invalidation algorithm that simply did not require so many auxiliary nodes would be a superior solution.
@@ 4. Spineless Traversal
Spineless Traversal improves on Double Dirty Bit by jumping directly between dirty nodes without accessing any auxiliary nodes; as a result, it suffers dramatically fewer cache misses. Achieving this requires a more computationally heavy approach: storing all dirty nodes in a priority queue and maintaining the correct traversal order using an order maintenance data structure. Spineless Traversal’s savings in cache misses typically outweigh the greater computational requirements of these data structures. Since Spineless Traversal is complex, this section develops it incrementally, first introducing the idea of a queue storing dirty nodes, then adding timestamps to maintain traversal order, and finally introducing the order maintenance structure to handle node insertion and deletion. A final step optimizes bulk insertions and deletions.
@@ 4.1. Jumping Directly to Dirty Nodes
To jump directly to dirty nodes, we introduce a queue that stores pointers to dirty nodes. More specifically, elements of the queue represent dirty bits in the layout tree that need to be cleared, and are represented in the queue by a pointer to the relevant layout node and an enumeration identifying which dirty bit needs to be cleared. When a dirty bit is set, the corresponding node and enumeration are added to the queue; the order of nodes in the queue will be explained later. Note that dirty bits are still present on layout nodes; elements are only added to the queue when a dirty bit is set for the first time, so the queue does not contain duplicates. To perform an incremental layout, elements are popped from the queue and the relevant fields on the relevant node are recomputed, and clearing the relevant dirty bit. In the process, new dirty bits may be set, so new elements may be added to the queue. This is repeated until the queue is empty. Only dirty nodes are ever in the queue, meaning only they—not auxiliary nodes—are ever accessed.
Concretely, the queue is stored in a packed array, meaning that clearing a dirty bit accesses one dirty element, plus its at most five neighbors. If the queue stays in L2 cache due to repeated access, while elements do not due to their size, this means clearing a dirty bit incurs a maximum of six L2 cache misses—much fewer than the number of auxiliary accesses typical of Double Dirty Bit.
@@ 4.2. Maintaining Queue Order
To ensure that each field is only recomputed once, queue elements must be in the right order. We thus add a timestamp field for every dirty bit on every node, giving timestamps the same relative order as in a from-scratch layout. In the full Spineless Traversal algorithm, as explained below, this timestamp is an “order maintenance object”, but for now the reader can imagine it an integer counting up from 0. The queue of dirty nodes is then refined to a priority queue, ordered by these timestamps. The priority queue “pops” the element with the lowest timestamp, thus ensuring that Spineless Traversal clears dirty bits in timestamp order and thus clears each dirty bit exactly once.
Concretely, we use a min-heap as our priority queue, which is cache-friendly and requires relatively few operations for each push and pop. Timestamps are stored adjacent to dirty bits, meaning they do not introduce any new L2 cache misses. The priority queue is typically small: while there are typically thousands of nodes, with each node having dozens of fields in our evaluation, the priority queue typically contains less than 1000 elements, and for the most latency-critical interactions, like hovers or drags, it can contain 100 or fewer. With such a small size, a priority queue push/pop requires 5–10 timestamp comparisons, which can be performed in roughly the time of one to three L2 cache misses in our optimized implementation.
@@ 4.3. Order Maintenance
The final challenge is efficiently assigning timestamps to every dirty bit on every node. Simple incrementing integer timestamps, sadly, don’t work when nodes are inserted or deleted: inserting a node would shift all future timestamps, so while inserting the node and performing incremental layout might access only a few nodes, reassigning timestamps would access many more. Instead, following SAC (Acar, 2005), Spineless Traversal uses an order maintenance data structure (OM) to assign timestamps. First introduced by Dietz (1982), order maintenance is a data structure that maintains a totally ordered set of objects while allowing objects to be added and removed from the order arbitrarily. Crucially, adding, removing, and comparing nodes takes $O(1)$ time. Abstractly, order maintenance provides the following API:
@@ 3. $\mathsf{Create}(p)$
Deleting OM objects is also possible, though by default our implementation does not do so. [^6] On long-running pages this is a memory leak, albeit a very slow one; enabling OM object deletion adds only a minor slowdown to Spineless Traversal, roughly 2%.
Our implementation is based on that by Bender et al. (2002), which uses a two-level structure with a double-linked list of double-linked lists. Objects are represented by nodes in the lower-level lists. Both levels are ordered; the total order traverses lower-level lists in order, in the order dictated by the higher-level list. Each object (node in the lower-level list) maintains a pointer to its higher-level list cell; two objects are in the same low-level list if they have the same higher-level pointer. To allow fast comparisons between nodes, both low-level and high-level list cells store an unsigned integer of fixed size (in our implementation, 32 bits) called labels. Within both lists, node labels are strictly increasing; this makes comparisons fast. Specifically, comparison has two cases: if the two objects are in the same low-level list, their labels are compared directly, while if they are in in different ones, their parents’ labels are compared. This comparison operation is the bulk of the Spineless Traversal time so its speed is essential; Section 6 discusses critical micro-optimizations that bring its latency down to around 5 cycles.
To create an object inside an order maintenance structure, a new lower-level list cell is created whose label is the average of the two neighboring labels.777 When creating a node after the last node, the maximum representable number is used as the larger number. If the two labels differ by exactly 1, however, this would repeat a label. In this case, the data structure re-balances itself, evenly reassigning labels to existing objects. This process might create a new higher-level list cell to split a lower-level list in two, ensuring a sufficiently large gap between its cells. Rebalancing is algorithmically tricky but is not a significant time sink in our use case, so we do not detail re-balancing here; details can be found in Bender et al. (2002).
@@
Concretely, the timestamp for each dirty bit is now represented by a pointer to the lower-level OM object. Priority queue elements are a node pointer, an enumeration naming the dirty bit, and a pointer to the OM node, padded to a total of 16 bytes to align with cache lines. To compare two timestamps, both pointers must be dereferenced to access the two lower-level OM objects. The associated higher-level OM objects must also be accessed. In the typical case, where the priority queue remains small, all of these OM objects in the priority queue are already in cache, so “pops” have no additional cache misses, while “pushes” add one, for the new element OM cell. In total, all the priority queue and order maintenance operations thus add the equivalent of 3–5 L2 cache misses in latency, about 100–200 cycles. Since recomputing a field, accessing neighbors’ fields, and setting neighbors’ dirty bits can itself take hundreds of cycles, Spineless Traversal’s overhead is fairly small.
@@ 4.4. Subtree Insertion
Bulk insertions into the layout tree are common in “lazy loading” patterns: a “shell” web page loads first and shows a loading indicator; then the “content” loads and is inserted a large subtree, replacing the loading indicator. This pattern is encouraged by frameworks like React, and can occur in several stages, with a “shell” first inserting “subshells” which themselves load subcomponents in turn. Efficiently handling these bulk insertions requires special care in Spineless Traversal, as bulk insertions are typically responsible for Spineless Traversal’s worst performance relative to Double Dirty Bit.
The basic issue is that in a fresh subtree being inserted into the page, every field needs to be computed and every OM object needs to be initialized, without causing the priority queue itself to grow large. Our solution adds these “initialization passes” as special elements in the priority queue: besides $(v,n)$ pairs for a dirty field $v$ on node $n$, an element can also be $(p,r)$, a pass $p$ that needs to initialize the entire subtree rooted at $r$. The timestamp for such a special element is just after the last field assigned by $p$ in $T$’s previous sibling or parent.888 An edge case is inserting a subtree into a subtree that itself has not yet been laid out. In this case no further actions need to be taken, since both subtrees will be visited together.
@@ 5.1. Compiler Interface
The Megatron compiler parses layout rules and schedules, and compiles them to fast incremental layout algorithms in C++. Concretely, output of Megatron is an recompute function, which is passed a node and field name and recomputes that field on that node, setting dirty bits for any dependent fields. To produce this output, the compiler analyzes the static dependencies of each field assignment, synthesizes dirty bit propagation code, and generates packed, cache-friendly data structures.
@@
To allow for a head-to-head comparison between Double Dirty Bit and Spineless Traversal, the driver program is parameterized over a simple “invalidation traversal” API. It has two key methods: a dirty method that marks a node’s field for later re-computation and a clean function that invokes recompute on all dirty fields. Signatures for both functions are shown in Figure 9; there are also boilerplate methods to initialize data structures. The dirty method is called both when modifying the layout tree and also from the recompute function itself, to dirty dependents when a field’s value changes. Three invalidation traversals are available in Megatron: a naive traversal that visits each node, Double Dirty Bit, and Spineless Traversal.

### v5

@@ 3.1. Dirty bits
Incremental layout is conceptually straightforward. Each layout field has a corresponding dirty bit, which defines whether that layout field needs to be recomputed. [^5] Packed layout fields can all have the same corresponding dirty bit, which is set if any of those fields need to be recomputed. APIs that write to an attribute or property, or add or remove layout nodes, set the dirty bits for all layout fields that read from those fields or node pointers. For example, in the paragraph layout of Figure 5, when a node’s height attribute is changed, its $H$ field must be marked dirty. When a subtree is deleted, its next sibling’s $HA$ field must be marked dirty. If the subtree was the last child of its parent, the parent’s $H$ field must also be marked dirty. This set of fields can be statically determined in our DSL and in fact our Megatron compiler synthesizes this marking code automatically.
Incremental layout then has the task of clearing dirty bits by recomputing dirty fields. Conceptually this is simple: find all dirty bits on all nodes and recompute the fields they correspond to. However, when a layout field is recomputed, its value might change, and then any layout fields that read the old value are would be out of date. Thus, when incremental layout modifies a layout field, it must dirty all layout fields that depend on it, “propagating” dirty bits through the document. For example, in the paragraph layout of Figure 5, when a node’s $H$ field is recomputed, its $HA$ field must be marked dirty. When that $HA$ field is recomputed, its next sibling’s $HA$ field is then marked dirty in turn. Because of dirty bit propagation, the set of layout fields currently marked dirty is distinct from the set of all layout fields that incremental layout must ultimately recompute.
This second set is discovered in the process of performing incremental layout: dirty bits are only propagated when a layout field’s value changes, which is only known when that field is recomputed and the old and new value can be compared. For example, suppose the user is typing out a paragraph of text, laid out using Figure 5’s layout algorithm. Most edits just add text to a single line, not affecting its height or available width; these only end up dirtying a single field on a single node. This is why incremental layout works: many interactions—especially latency-critical ones like hovers, drags, animations, and text editing—do not change fields on containing elements, and end up affecting only a small portion of the page.
@@ 3.2. Recomputation Order
Incremental layout strives to minimize the number of layout fields recomputed. That requires ensuring that, once recomputed, a layout field is not marked dirty again during that layout. Luckily, there is a simple way to guarantee this: a correct from-scratch layout always processes a field’s dependencies before the field itself, so recomputing layout fields in the same relative order as a from-scratch layout ensures that each layout field is marked dirty at most once. A naive incremental layout algorithm might thus execute the same layout schedule as a from-scratch layout but skip re-computing any fields that aren’t dirty. This algorithm recomputes the minimal number of layout fields, clears all the dirty bits, and avoids ever recomputing a layout field more than once.
However, this naive algorithm isn’t particularly fast: it must access every layout node to check its dirty bits, which incurs a lot of cache misses. In the common case, most dirty bits aren’t set; these “auxiliary” accesses cause needless cache misses and end up wasting a significant fraction of run time. For incremental layout to be fast, auxiliary accesses—accesses that don’t result in field recomputations—must be minimized.
@@ 3.3. The Double Dirty Bit Algorithm
The state-of-the-art Double Dirty Bit algorithm uses summary bits to reduce auxiliary accesses; it is used, with variations, in all major rendering engines. For each dirty bit, we add a second “summary bit”, which is set for any node where the corresponding dirty bit is set anywhere in the subtree rooted at that node. When a field is marked dirty, its dirty bit is set, and the associated summary bit is also set on that node and all its ancestors. When performing incremental layout, subtrees whose summary bits are clear can be skipped. Figure 7 contains pseudo-code for Double Dirty Bit as an iterator over dirty nodes.
@@
    self.dirty_bit = True
@@
        self.dirty_bit = False
@@
Skipping subtrees with no dirty bits reduces the number of auxiliary accesses and greatly improves performance (Garseil, [n. d.]; Panchekha and Harrelson, 2024). However, some auxiliary accesses remain. Specifically, when a dirty bit is set on some node, summary bits will be set for all ancestors of that node, which Double Dirty Bit will have to check. Moreover, for each ancestor, Double Dirty Bit will recurse into its children, adding even more auxiliary accesses. Since layout trees are poorly balanced and often very wide, the number of auxiliary accesses can be large. For example, a single dirtied node in Figure 1 has 66 auxiliary nodes, and on larger pages the dirty-to-auxiliary ratio can be even worse.
Developers can sometimes reduce the number of auxiliary accesses by reorganizing large, complex pages; performance monitoring tools like Google’s Lighthouse (Team, [n. d.]) will suggest doing so. But this can require global changes to the shape of the layout tree, which (in modern frameworks like React) requires refactoring the application as a whole. Naturally, an invalidation algorithm that simply did not require so many auxiliary nodes would be a superior solution. And while this is a quite specialized performance problem, it is a major source of latency in existing web browsers, which in turn are both critical application platforms and also already highly optimized, meaning remaining sources of latency are particularly challenging.
@@ 4. Spineless Traversal
Spineless Traversal improves on Double Dirty Bit by jumping directly between dirty nodes without accessing any auxiliary nodes; as a result, it suffers dramatically fewer cache misses. Achieving this requires a more computationally heavy approach: storing all dirty nodes in a priority queue and maintaining the correct traversal order using an order maintenance data structure. Spineless Traversal’s savings in cache misses typically outweigh the greater computational requirements of these data structures. Since Spineless Traversal is complex, this section develops it incrementally, first introducing the idea of a queue storing dirty nodes, then adding timestamps to maintain traversal order, and finally introducing the order maintenance structure to handle node insertion and deletion. A final subsection optimizes bulk insertions and deletions.
@@ 4.1. Jumping Directly to Dirty Nodes
To jump directly to dirty nodes, we introduce a queue of pointers to dirty nodes. More specifically, elements of the queue represent dirty bits in the layout tree that need to be cleared, and are represented in the queue by a pointer to the relevant layout node and an enumeration identifying which dirty bit needs to be cleared. When a dirty bit is set, the corresponding node and enumeration are added to the queue; the order of nodes in the queue will be explained later. Note that dirty bits are still present on layout nodes; elements are only added to the queue when a dirty bit is set for the first time, so the queue does not contain duplicates. To perform an incremental layout, elements are popped from the queue and the relevant fields on the relevant node are recomputed, clearing the relevant dirty bit. In the process, new dirty bits may be set, so new elements may be added to the queue. This is repeated until the queue is empty. Only dirty nodes are ever in the queue, meaning only they—not auxiliary nodes—are ever accessed.
Concretely, the queue is stored in a packed array, which due to heavy use will remain in cache. Clearing a dirty bit thus suffers cache misses only for the dirty element, plus its at most five neighbors. This means clearing a dirty bit incurs a maximum of six L2 cache misses—much fewer than the number of auxiliary accesses typical of Double Dirty Bit.
@@ 4.2. Maintaining Queue Order
To ensure that each field is only recomputed once, queue elements must be in the right order. We thus add a timestamp field for every dirty bit on every node, giving timestamps the same relative order as in a from-scratch layout. In the full Spineless Traversal algorithm, as explained below, this timestamp is an “order maintenance object”, but for now the reader can imagine it an integer counting up from 0. The queue of dirty nodes is then refined to a priority queue, ordered by these timestamps. The priority queue “pops” the lowest timestamp first, so Spineless Traversal clears dirty bits in timestamp order and thus clears each dirty bit exactly once.
Concretely, we use a min-heap as our priority queue, which is cache-friendly and requires relatively few operations for each push and pop. Timestamps are stored adjacent to dirty bits, meaning they do not introduce any new L2 cache misses. The priority queue is typically small: while a web page may have thousands of nodes, with each node having dozens of fields in our evaluation, the priority queue typically contains less than 1000 elements, and for the most latency-critical interactions, like hovers or drags, it can contain 100 or fewer. With such a small size, a priority queue push/pop requires 5–10 timestamp comparisons, which can be performed in roughly the time of one to three L2 cache misses in our optimized implementation.
@@ 4.3. Order Maintenance
The final challenge is efficiently assigning timestamps to every dirty bit on every node. Simple incrementing integer timestamps, sadly, don’t work: inserting a node would requiring adjusting all later timestamps, which would introduce its own set of auxiliary accesses. Instead, following SAC (Acar, 2005), Spineless Traversal uses an order maintenance data structure (OM) to assign timestamps. First introduced by Dietz (1982), order maintenance maintains a totally ordered set of objects while allowing objects to be added and removed from the order arbitrarily. Crucially, adding, removing, and comparing nodes takes $O(1)$ time. Abstractly, order maintenance provides the following API:
@@ 3. $\mathsf{Create}(p)$
Deleting OM objects is also possible, though by default our implementation does not do so. [^6] On long-running pages this causes very-slowly-growing memory usage. Enabling OM object deletion avoids this but adds roughly 2% to Spineless Traversal’s running time.
Our implementation is based on that by Bender et al. (2002), which uses a two-level structure with a doubly-linked list of doubly-linked lists (Figure 8). Objects are represented by nodes in the lower-level lists. Both levels are ordered; the total order is lexicographic, traversing higher-level lists in order and then, for each higher-level list, its lower-level list in order. Each object (node in the lower-level list) maintains a pointer to its higher-level list cell; two objects are in the same low-level list if they have the same higher-level pointer. To allow fast comparisons between nodes, each low-level and high-level list cells stores an unsigned integer of fixed size (in our implementation, 32 bits) called its label. Within a list, node labels are strictly increasing; this makes comparisons fast. Specifically, comparison has two cases: if the two objects are in the same low-level list, their labels are compared directly, while if they are in in different ones, their parents’ labels are compared. This comparison operation is the bulk of the Spineless Traversal time so its speed is essential; Section 6 discusses critical micro-optimizations that bring its latency down to around 5 cycles.
To create an object inside an order maintenance structure, a new lower-level list cell is created whose label is the average of the two neighboring labels.777 When creating a node after the last node, the maximum representable number is used as the larger number. If the two labels differ by exactly 1, however, this would repeat a label. In this case, the data structure re-balances itself, evenly reassigning labels to existing objects. This process might create a new higher-level list cell to split a lower-level list in two, ensuring a sufficiently large gap between its cells. Rebalancing is algorithmically tricky but is not a significant time sink in our use case, so we do not detail re-balancing here; interested readers can find a description in Bender et al. (2002).
@@
Concretely, the timestamp for each dirty bit is now represented by a pointer to the lower-level OM object. Priority queue elements are a node pointer, an enumeration naming the dirty bit, and a pointer to the OM node, padded to a total of 16 bytes to align with cache lines. To compare two timestamps, both pointers must be dereferenced to access the two lower-level OM objects. The associated higher-level OM objects must also be accessed. In the typical case, where the priority queue remains small, all of the OM objects in the priority queue are already in cache, so “pops” have no additional cache misses, while “pushes” add just one or two, for the new element OM cells. In total, all the priority queue and order maintenance operations thus add the equivalent of 3–6 L2 cache misses in latency, about 100–300 cycles. Since recomputing a field, accessing neighbors’ fields, and setting neighbors’ dirty bits can itself take hundreds of cycles, Spineless Traversal’s overhead is fairly small.
@@ 4.4. Subtree Insertion
Bulk insertions into the layout tree are common in “lazy loading” patterns: a “shell” web page loads first and shows a loading indicator; then the “content” loads and is inserted as a large subtree, replacing the loading indicator. This pattern is encouraged by frameworks like React, and can occur in several stages, with a “shell” first inserting “subshells” which themselves load subcomponents in turn. Efficiently handling these bulk insertions requires special care in Spineless Traversal, as bulk insertions are typically responsible for Spineless Traversal’s worst performance relative to Double Dirty Bit.
The basic issue is that when a fresh subtree is inserted, every field needs to be computed and every OM object needs to be initialized, all without causing the priority queue to grow too large. Our solution add is to add “initialization passes” as special elements in the priority queue: besides $(v,n)$ pairs for a dirty field $v$ on node $n$, an element can also be $(p,r)$, a pass $p$ that needs to initialize the entire subtree rooted at $r$. The timestamp for such a special element is just after the last field assigned by $p$ in $T$’s previous sibling or parent.888 An edge case is inserting a subtree into a subtree that itself has not yet been laid out. In this case no further actions need to be taken, since both subtrees will be visited together.
@@ 5.1. Compiler Interface
The Megatron compiler parses layout rules and schedules and compiles them to a fast incremental C++ layout algorithm. Concretely, the output of Megatron is a recompute function, which is passed a node and a field name and which recomputes that field on that node, setting dirty bits for any dependent fields. To produce this output, the compiler analyzes the static dependencies of each field assignment, synthesizes dirty bit propagation code, and generates packed, cache-friendly data structures.
@@
To allow for a head-to-head comparison between Double Dirty Bit and Spineless Traversal, the driver program is parameterized over a simple “invalidation traversal” API. It has two key methods: a dirty method that marks a node’s field for later re-computation and a clean method that invokes recompute on all dirty fields. Signatures for both functions are shown in Figure 9; there are also boilerplate methods to initialize data structures. The dirty method is called both when modifying the layout tree and also from the recompute function itself, to dirty dependents when a field’s value changes. Three invalidation traversals are available in Megatron: a naive traversal that visits each node, Double Dirty Bit, and Spineless Traversal.